OpenStudy (anonymous):

A woman is 66 7/12" tall and her son is 58 7/9" tall. How many inches taller is the woman?

OpenStudy (anonymous):

\[ 66\frac{7}{12} -58\frac{7}{9} \] \[\frac{799}{12}-\frac{529}{9} \]

OpenStudy (anonymous):

Then you can get the numerator to be the same, and find the difference of the fraction to get the difference of heights.

OpenStudy (anonymous):

My calculator died on me, so you'll have to figure that bit out on your own.

OpenStudy (anonymous):

not sure of next step?

OpenStudy (anonymous):

Ok, well, this'll take some time, but I'll show you.

OpenStudy (anonymous):

I thought answer was 12".

OpenStudy (anonymous):

\[\frac{799\times12}{108}- \frac{529\times12 }{108} \]

OpenStudy (anonymous):

sorry, 799times9

OpenStudy (anonymous):

how do you get the 108?

OpenStudy (anonymous):

ok. 9 x12?

OpenStudy (anonymous):

you do 799 times 9 and 529 times 12, so that the bottom can equal 108

OpenStudy (anonymous):

\[ \frac{7191-6348}{108} \]

OpenStudy (anonymous):

\[ \frac{843}{108}inches\] should be the answer which you can simplify

OpenStudy (anonymous):

Am so lost!

OpenStudy (anonymous):

Where did you get lost?

OpenStudy (anonymous):

how did you get to the difference in inches when the fractions are so uneven?

OpenStudy (anonymous):

I made the fractions even by making them first into Whole fractions, and then making the denominators the same so that I could take away both fractions. Then I got a large Whole fraction, which I placed again into a simplified verion, which is actually 7"87/108 now that I see it.

OpenStudy (anonymous):

so 66 7/12 = p (mother) 58 7/9= h (son) and subtract h from p to get ?

OpenStudy (anonymous):

to get the difference in height of the two, which results in how much taller the woman is than the son.

OpenStudy (anonymous):

p= 799*9=7191/108 and h= 529*12=6348/108 reduce to 843/108?

OpenStudy (anonymous):

Yup, which is the answer. You can get that as a proper fraction by dividing 843 by 108 and finding the remainder, then placing the remainder over 108, which gives 7 87/108"

OpenStudy (anonymous):

so the mother is 7" taller?

OpenStudy (anonymous):

yup, 7 and 87/108" taller

OpenStudy (anonymous):

or does it have to be 7 87/108" taller?

OpenStudy (anonymous):

Yes, it has to be the second as it's 7.80 inches if you use it as a decimal point.

OpenStudy (anonymous):

7 87/108" taller

OpenStudy (anonymous):

how do i simplify 7 87/108?

OpenStudy (anonymous):

7 29/36

OpenStudy (anonymous):

that's the most simplified verion.

OpenStudy (anonymous):

so the lowest divisible is 3 right?

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

THANK YOU!

OpenStudy (anonymous):

you're welcome :) You understand it now?

OpenStudy (anonymous):

yes thanks.

OpenStudy (anonymous):

great :)